# Proximality and selflessness for group C*-algebras

### v1

@@ Abstract.
We prove that the reduced group $\mathrm{C}^{*}$-algebra of an infinite discrete group having a topologically-free extreme boundary is selfless in the sense of L. Robert.
@@ 1. Introduction
A remarkable property called selflessness is recently introduced for a $\mathrm{C}^{*}$-algebra by L. Robert ([Ro]) and it quickly attracted a number of researchers’ attention (see, e.g., [AGKEP], [HKER], [RTV], [Vi]) as it implies many important regularity properties such as simplicity, (in the tracial setting) stable rank one, and strict comparison. Let’s say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if its reduced group $\mathrm{C}^{*}$-algebra is selfless. In this paper, we prove that every infinite countable discrete group with a minimal topologically-free extremely-proximal action is $\mathrm{C}^{*}$-selfless. Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XC in [STW]). Our result should be compared with Kalantar and Kennedy’s theorem ([KK]) that a discrete group $\Gamma$ is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action. As the names suggest, strong proximality is weaker than extreme proximality ([Gl]).
@@
For $x\in X$, we denote by $\Gamma_{x}\coloneq\{g\in\Gamma:gx=x\}$ the stabilizer subgroup at $x$. The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer group is dense in $X$. Note that if an extreme boundary $\Gamma\curvearrowright X$ is not topologically-free, then every proper closed subset of $X$ is pointwise fixed by some $g\in\Gamma\setminus\{1\}$.
We say a sequence $(z_{n})_{n}\in\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that every topologically-free extreme boundary $X$ with $|X|>2$ accommodates an axial sequence.
@@
An infinite countable discrete group $\Gamma$ having a topologically-free extreme boundary is $\mathrm{C}^{*}$-selfless. More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$, the homomorphism
@@
given by the constant embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[\lambda(z_{n})]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
See, e.g., [Bo], [BIO], [FLMMS], [IO], [JR], [LS], [LBMB] for examples of extreme boundaries. Every group satisfying the conclusion of Theorem 1 is mixed-identity-free ([HO]) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary. In particular, non-degenerate free product groups are $\mathrm{C}^{*}$-selfless. For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for the characterization of topological freeness of the corresponding Bass–Serre tree boundaries. Non-elementarily relatively hyperbolic groups ([Bo]) with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications are topologically-free extreme boundaries. It is not clear whether this generalizes to acylindrically hyperbolic groups, because their hyperbolic boundaries need not be compact.
@@
A sequence is written as $(x_{n})_{n}$, while the equivalence class it represents in the ultrapower is written by $x_{n}]_{n}$. We use the symbol $z^{\pm}$ for a slightly abusive notation for “$z$ and/or $z^{-1}$” and similarly for $z_{\pm}$. For $\epsilon\in\{\pm\}$, we denote by $\bar{\epsilon}$ the opposite of $\epsilon$.
@@ Acknowledgments
This research was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 4–8, 2025. He acknowledges the kind hospitality and the exciting environment provided by the institute. He thanks G. Patchell for the stimulating talk there and a useful comment on this paper.
@@ 2. Trees and covariant representations
We recall the notion of trees and their compactification. A tree is a connected graph without nontrivial cycles. See Section 5.2 and Appendix E in [BO] for more on this theme. Here we consider a countable tree $\mathrm{T}$ of infinite degree. The tree $\mathrm{T}$ is identified with its vertex set. A geodesic ray is an infinite sequence $\omega(0),\omega(1),\ldots$ in $\mathrm{T}$ such that $\omega(n)$ adjacent to $\omega(n-1)$ and $\omega(n)\neq\omega(n-2)$. Geodesic rays $\omega$ and $\omega^{\prime}$ are equivalent if there are $N$ and $N^{\prime}$ such that $\omega(N+n)=\omega^{\prime}(N^{\prime}+n)$ for all $n\geq 0$. The boundary $\partial\mathrm{T}$ is the equivalence classes of geodesic rays and $\bar{\mathrm{T}}\coloneq\mathrm{T}\cup\partial\mathrm{T}$ is the compactification of $\mathrm{T}$ (whose topology will be introduced shortly). We define $f\colon\mathrm{T}\times\bar{\mathrm{T}}\to\mathrm{T}$ as follows. For $s\in\mathrm{T}$, we set $f(s,s)=s$. For distinct $s,t\in\mathrm{T}$, we define $f(t,s)\in\mathrm{T}$ to be the unique point between $t$ and $s$ that is adjacent to $t$. For $t\in\mathrm{T}$ and $\omega\in\partial\mathrm{T}$, we define $f(t,\omega)\coloneq f(t,\omega(n))$ for any $n$ large enough. For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$. That is, every sequence in $\mathrm{T}$ that diverges to infinity converges to $t$ in $\mathrm{T}_{t}$. We consider the embedding
@@
and equip $\bar{\mathrm{T}}$ with the induced topology. This makes $\bar{\mathrm{T}}$ a second-countable compact topological space (Proposition 5.2.5 in [BO]). Note that if $s\in\mathrm{T}$ and $t_{n}\in\mathrm{T}$ are such that $f(s,t_{n})\to\infty$ in $\mathrm{T}$, then $t_{n}\to s$ in $\bar{\mathrm{T}}$.
@@
Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a unital $\mathrm{C}^{*}$-algebra $A$. Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be a covariant representation. Assume that there is a $\Lambda$-equivariant unital positive map from $C(\bar{\mathrm{T}})$ into the center of $A$ and that for every $t]\in\mathrm{T}/\Lambda$, the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}A$. Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}A$
@@
Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be given. We denote by $\lambda$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $A$. We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\lambda\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in A$. Indeed, there is a sequence $\zeta_{n}\colon\bar{\mathrm{T}}\to\mathrm{Prob}(\mathrm{T})$ of Borel maps such that $\sup_{x\in\bar{\mathrm{T}}}\|g\zeta_{n}^{x}-\zeta_{n}^{gx}\|_{1}\to 0$ for every $g\in\Lambda$ (Lemma 5.2.6 in [BO]). It follows that $h_{n}^{t}(x)\coloneq\zeta_{n}^{x}(t)$ defines sequences of positive Borel functions on $\bar{\mathrm{T}}$ such that $\sum_{t\in\mathrm{T}}h_{n}^{t}=1$ and
@@
NB: the above sum is an infinite sum of positive Borel functions. Let $\theta$ denote the $\Lambda$-equivariant unital completely positive map from $C(\bar{\mathrm{T}})$ into the center of $A$. The map $\pi\circ\theta$ canonically extends to a normal unital positive map $\Theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of the bounded Borel functions into $(\pi\circ\theta)(C(\bar{\mathrm{T}}))^{\prime\prime}$. By “normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence. We note that $\Theta$ satisfies $|\Theta(f)|\leq\Theta(|f|)$ because the range of $\Theta$ is commutative. We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\Theta(h_{n}^{t})^{1/2}\xi$, or $V_{n}$ is the column operator $(\Theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$. Then
$\|V_{n}\pi(g)-(\lambda\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\Theta(h_{n}^{gt})^{1/2}-\Theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
@@
for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in A$. This proves the claim that $\pi$ is weakly contained in $\tilde{\pi}$. Thus continuity of $\pi$ follows from that of $\tilde{\pi}$.
@@
For the continuity of $\tilde{\pi}$, it suffices to consider the continuity of the vector states $\varphi$ associated with the vectors of the form $\delta_{t}\otimes\xi$, since they form a cyclic family. But $\varphi=\psi\circ E$, where $E$ is the conditional expectation from $\Lambda\ltimes_{\max}A$ onto $\Lambda_{t}\ltimes_{\max}A$ and $\psi$ is the vector state associated with $\xi$. (We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}A\subset\Lambda\ltimes_{\max}A$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}A$ follows from that for $\Lambda_{t}\ltimes_{\min}A$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit. ∎
3. A space with “tree structure”

### v6

@@ Abstract.
We prove that the reduced group $\mathrm{C}^{*}$-algebras of infinite countable discrete groups having topologically-free extreme boundaries, or more generally groups that satisfy certain combinatorial property including all acylindrically hyperbolic groups with no nontrivial finite normal subgroups and all Zariski-dense subgroups of $\mathrm{PSL}(n,\mathbb{R})$, are selfless in the sense of L. Robert. This generalizes the recent results of Amrutam, Gao, Kunnawalkam Elayavalli, and Patchell, and of Vigdorovich. We also prove that selflessness is stable under tensor product among exact $\mathrm{C}^{*}$-algebras and that a $\mathrm{C}^{*}$-probability space is selfless provided that it is either simple and purely infinite or simple, exact, $\mathcal{Z}$-stable, and uniquely tracial.
@@ 1. Introduction
A remarkable property called selflessness is recently introduced for a $\mathrm{C}^{*}$-algebra by L. Robert ([Ro]) and it quickly attracted a number of researchers’ attention (see, e.g., [AGKEP], [HKER], [RTV], [Vi]) as it implies many important regularity properties such as simplicity, (in the tracial setting) stable rank one, and strict comparison. In particular, Amrutam, Gao, Kunnawalkam Elayavalli, and Patchell ([AGKEP]) has proved that a large family of groups, namely acylindrically hyperbolic groups with the rapid decay property and with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless. Here, we say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if it its reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(\Gamma)$ is selfless.
In this paper, we prove that every infinite countable discrete group with a minimal topologically-free extremely-proximal action is $\mathrm{C}^{*}$-selfless. Our method is topological, as opposed to the more analytical one in [AGKEP], and to construct a kind of “tree-graded” space ([DS]) out of an extreme boundary (see Section 3). Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XCI in [STW]). We remind Kalantar and Kennedy’s theorem ([KK]) that a discrete group is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action. As the names suggest, strong proximality is weaker than extreme proximality ([Gl]).
@@
For $x\in X$, we denote by $\Gamma_{x}\coloneq\{g\in\Gamma:gx=x\}$ the stabilizer subgroup at $x$. The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer groups is dense in $X$. Note that if an extreme boundary $\Gamma\curvearrowright X$ is not topologically-free, then every proper closed subset of $X$ is pointwise fixed by some $g\in\Gamma\setminus\{1\}$.
We say a sequence $(z_{n})_{n}$ in $\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that an extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
@@
An infinite countable discrete group $\Gamma$ having a topologically-free extreme boundary is $\mathrm{C}^{*}$-selfless. More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$ on $\mathbb{N}$, the homomorphism
@@
given by the diagonal embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[z_{n}]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
See, e.g., [Bo], [BIO], [FLMMS], [IO], [JR], [LS], [LBMB] for examples of extreme boundaries. Every group satisfying the conclusion of Theorem 1 is mixed-identity-free (see [HO] for the terminology) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary. In particular, non-elementary free product groups are $\mathrm{C}^{*}$-selfless. For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for a characterization of topological freeness of the corresponding Bass–Serre tree compactifications. Non-elementarily relatively hyperbolic groups with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications ([Bo]) are topologically-free extreme boundaries. It is not clear whether this generalizes to acylindrically hyperbolic groups (with no assumption on the rapid decay property, cf. [AGKEP]), because their hyperbolic structure need not yield suitable compactifications. We fix this problem in Section 8 and prove that a group satisfying property $\mathrm{P}_{\mathrm{PHP}}$ (defined there) is $\mathrm{C}^{*}$-selfless. The class of groups with property $\mathrm{P}_{\mathrm{PHP}}$ contains all acylindrically hyperbolic groups with no nontrivial finite normal subgroups as well as all Zariski-dense subgroups of $\mathrm{PSL}(d\geq 2,\mathbb{R})$ (cf. [Vi]).
There are $\mathrm{C}^{*}$-selfless groups outside the mixed-identity-free realm.
Theorem 2.
Let $(A_{i},\varphi_{i})$ be separable $\mathrm{C}^{*}$-probability spaces. Assume that all $(A_{i},\varphi_{i})$ are selfless and exact. Then the tensor product $\bigotimes_{i}(A_{i},\varphi_{i})$ is selfless. In particular, the class of countable discrete groups that are $\mathrm{C}^{*}$-selfless and exact is closed under direct product.
We introduce the notion of complete selflessness in Section 6 and prove that the examples in Theorem 1 and their tensor product are in fact completely selfless, while no nuclear tracial $\mathrm{C}^{*}$-probability space is completely selfless. The following is a partial converse to Theorem 3.1 in [Ro] and solves Question 5.4 in [Ro]. Note that by Matui and Sato’s theorem ([MS]), $\mathcal{Z}$-stability is necessary for a nuclear tracial $\mathrm{C}^{*}$-probability space to be selfless.
Theorem 3.
A simple and purely infinite $\mathrm{C}^{*}$-probability space $(A,\varphi)$ is completely selfless. A simple, exact, $\mathcal{Z}$-stable, and uniquely tracial $\mathrm{C}^{*}$-probability space $(A,\tau)$ is selfless.
@@
We use the symbol $\pm$” for a slightly abusive notation for “++ and/or -” or “+1+1 and/or 1-1”. We denote by $A^{\mathcal{U}}$ the ultrapower of a $\mathrm{C}^{*}$-algebra $A$ and by $a_{n}]_{n}$ the element in $A^{\mathcal{U}}$ represented by a sequence $(a_{n})_{n}$ in $A$. A $\mathrm{C}^{*}$-probability space is a pair $(A,\varphi)$ of a unital $\mathrm{C}^{*}$-algebra $A$ and a distinguished state $\varphi$ which is not necessarily faithful but assumed GNS-faithful (a.k.a. non-degenerate) unless the $\mathrm{C}^{*}$-algebra in consideration is an ultrapower or an ultraproduct.
@@ Acknowledgments
A part of this work was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 3–8, 2025 and the Isaac Newton Institute for the program “Operators, Graphs, Groups” in October 2025. He acknowledges the kind hospitality and the exciting environment provided by the institutes. He thanks G. Patchell for the stimulating talk at MFO and a useful comment on this paper and E. Breuillard and I. Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$.
@@ 2. Trees and covariant representations
We recall the notion of trees and their compactification. A tree is a connected graph without nontrivial cycles. See Section 5.2 and Appendix E in [BO] for more on this theme. Here we consider a countable tree $\mathrm{T}$ (of infinite degree). The tree $\mathrm{T}$ is identified with its vertex set. A geodesic ray is an infinite sequence $\omega(0),\omega(1),\ldots$ in $\mathrm{T}$ such that $\omega(n)$ adjacent to $\omega(n-1)$ and $\omega(n)\neq\omega(n-2)$. Geodesic rays $\omega$ and $\omega^{\prime}$ are equivalent if there are $N$ and $N^{\prime}$ such that $\omega(N+n)=\omega^{\prime}(N^{\prime}+n)$ for all $n\geq 0$. The boundary $\partial\mathrm{T}$ is the equivalence classes of geodesic rays and $\bar{\mathrm{T}}\coloneq\mathrm{T}\cup\partial\mathrm{T}$ is the compactification of $\mathrm{T}$ (whose topology will be introduced shortly). We define $f\colon\mathrm{T}\times\bar{\mathrm{T}}\to\mathrm{T}$ as follows. For $s\in\mathrm{T}$, we set $f(s,s)=s$. For distinct $s,t\in\mathrm{T}$, we define $f(t,s)\in\mathrm{T}$ to be the unique point between $t$ and $s$ that is adjacent to $t$. For $t\in\mathrm{T}$ and $\omega\in\partial\mathrm{T}$, we define $f(t,\omega)\coloneq f(t,\omega(n))$ for any $n$ large enough. For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ (or the union of $t$ and its adjacent points) and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$. We consider the embedding
@@
and equip $\bar{\mathrm{T}}$ with the induced topology. This makes $\bar{\mathrm{T}}$ a second-countable compact topological space such that every automorphism on $\mathrm{T}$ extends to a homeomorphism on $\bar{\mathrm{T}}$ (Proposition 5.2.5 in [BO]). Note that if $s\in\mathrm{T}$ and $t_{n}\in\mathrm{T}$ are such that $f(s,t_{n})\to\infty$ in $\mathrm{T}$, then $t_{n}\to s$ in $\bar{\mathrm{T}}$.
@@
Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a $\mathrm{C}^{*}$-algebra $C$. Let $\pi\colon\Lambda\ltimes_{\max}C\to\mathbb{B}(\mathcal{H})$ be a covariant representation and assume that there is a $\Lambda$-equivariant representation of $C(\bar{\mathrm{T}})$ into the commutant $\pi(C)^{\prime}$ of $\pi(C)$. Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}C$ if for every $t]\in\mathrm{T}/\Lambda$ the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}C$.
@@
Let $\pi\colon\Lambda\ltimes_{\max}C\to\mathbb{B}(\mathcal{H})$ be given. We denote by $\tau$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $G$. We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\tau\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in C$. Indeed, there is a sequence $\zeta_{n}\colon\bar{\mathrm{T}}\to\mathrm{Prob}(\mathrm{T})$ of Borel maps such that $\sup_{x\in\bar{\mathrm{T}}}\|g\zeta_{n}^{x}-\zeta_{n}^{gx}\|_{1}\to 0$ for every $g\in\Lambda$ (Lemma 5.2.6 in [BO]). It follows that $h_{n}^{t}(x)\coloneq\zeta_{n}^{x}(t)$ defines sequences of positive Borel functions on $\bar{\mathrm{T}}$ such that $\sum_{t\in\mathrm{T}}h_{n}^{t}=1$ and
@@
NB: the above sum is an infinite sum of positive Borel functions. The $\Lambda$-equivariant representation of $C(\bar{\mathrm{T}})$ into $\pi(C)^{\prime}$ uniquely extends to a $\sigma$-normal $\Lambda$-equivariant representation $\theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of bounded Borel functions into $\pi(C)^{\prime}$. By $\sigma$-normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence. We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\theta(h_{n}^{t})^{1/2}\xi$, or equivalently, $V_{n}$ is the column operator $(\theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$. Then
$\|V_{n}\pi(g)-(\tau\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\theta(h_{n}^{gt})^{1/2}-\theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
@@
for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in C$. This proves the claim that $\pi$ is weakly contained in $\tilde{\pi}$. Thus continuity of $\pi$ follows from that of $\tilde{\pi}$.
@@
For the continuity of $\tilde{\pi}$, it suffices to consider the continuity of the vector states $\varphi$ associated with the vectors of the form $\delta_{t}\otimes\xi$, since they form a cyclic family. But $\varphi=\psi\circ E$, where $E$ is the canonical conditional expectation from $\Lambda\ltimes_{\max}C$ onto $\Lambda_{t}\ltimes_{\max}C$ and $\psi$ is the vector state associated with $\xi$. (We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}C\subset\Lambda\ltimes_{\max}C$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}C$ follows from that for $\Lambda_{t}\ltimes_{\min}C$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit. ∎
3. Extreme boundaries and “tree-graded” spaces